Given 25.0 g of Al, how many grams of the product Al2O3 could be produced?
Answer #1

4 Al + 3 O2 = 2 Al2O3

or,   Al + (3/4) O2 = (2/4) Al2O3

25 gm Al = (25/27) = 0.93 moles of Al

1 mole Aluminum produces 0.5 mole of  Al2O3

0.93 moles of Aluminum produces (0.5 X 0.93) = 0.465 moles of Al2O3

Molar weight of Al2O3 = ( 27X2) + (16 X3) = 102

0.465 moles of Al2O3 = 0.465 X 102 = 47.43 gms of  Al2O3